# On the Design and Analysis of LLM-Based Algorithms

### v2

@@ 3.3.2 Analysis
[…]
@@
[…]
@@ 3.3.3 Experiments
Results with Llama-3-8B.
Our empirical results with a Llama-3-8B model are illustrated in Figure 4 and explained in the following.
In Figure 4(a), we vary the problem size $n$ and set $m=n$ in our algorithm, which means the algorithm becomes equivalent to a single LLM call. Unsurprising, error metrics in this counting task monotonely increase with $n$. The number of prefilling tokens increases linearly with $n$, while the number of decoding tokens is insensitive to $n$, since we prompt the LLM to output its answer directly without intermediate reasoning steps. The latency of one LLM call also increases linearly with $n$, which is likely due to the relatively small sequence lengths and the compute-bound nature of LLM inference by a Llama-3-8B model running on a CPU.
In Figure 4(b), we fix $n=200$ and vary the hyperparameter $m$, namely the size of each sub-task in our proposed LLM-based algorithm. It is confirmed that decomposing the original task into smaller parallel sub-tasks with a smaller value of $m$ improves the accuracy of the overall algorithm, while incuring higher cost metrics, except for the latency with infinite parallelism (which is monotonely increasing in $m$) and the latency with parallelism degree $p=4$ (which achieves the minimum at $m=n/p=50$, as was predicted by our previous analysis).
@@
Figure 5 demonstrates the empirical results for the same experiments but with a GPT-4-Turbo model. We observe from Figure 5(a) that the latency of one LLM call is insensitive to the input problem size $n$, which is likely because LLM inference of GPT-4-Turbo is memory-bound for the range of sequence lengths considered in our experiments. One potential implication is that, for the LLM-based algorithm with parallel decomposition, the latency with infinite parallelism $p=\infty$ might slightly increase for smaller $m$ and hence larger $k$, due to the random variation of latencies in reality; this is indeed what we observe from Figure 5(b). Other than that, the results in Figure 5 are similar to those in Figure 4.
@@
Figure 4: Empirical results for counting with Llama-3-8B.
@@
3.4 Example: sorting
For a more challenging example, let us consider the classical sorting problem: given a list of $n$ numbers $\bm{x}\in{\mathbb{R}}^{n}$, the task is to sort it in ascending order.
3.4.1 Algorithm
Below is an LLM-based algorithm for sorting a list, which generalizes the naive approach of sorting the list with one single LLM call:
  1. 1.
    Divide the input list $\bm{x}$ into $k$ disjoint sub-lists $\bm{x}_{1},\dots,\bm{x}_{k}$ of lengths $m_{1},\dots,m_{k}$;
  2. 2.
    For each $i\in[k]$, use one LLM call to sort the $i$-th sub-list, which returns a solution $\bm{y}_{i}$;
  3. 3.
    Merge the sub-lists $\bm{y}_{1},\dots,\bm{y}_{k}$ into a single list $\bm{y}$ using a symbolic algorithm.
We note two details about this algorithm. (1) To ensure efficiency and stability, for each LLM call, we prompt the LLM to generate the sorted list directly, without intermediate reasoning steps. It has been verified empirically that the LLMs considered in our experiments can follow such instructions, and generate text of length $\mathcal{L}_{\textsf{dec}}(m)=O(m)$ that can be easily parsed into a list. (2) Step 3 of the algorithm relies on a classical symbolic algorithm for merging two sorted lists, which maintains two moving pointers, one for each list, and chooses each entry of the merged list by comparing the values corresponding to the two pointers. Merging multiple sorted lists can be done by merging one pair of lists at a time, in an incremental or hierarchical manner. Python code for these procedures can be found in Listing 1 at the end of this subsection. Although they are designed under the assumption that the input lists are sorted, they can also be applied to input lists that are not fully sorted, which can possibly happen within the LLM-based algorithm since the input lists in Step 3 are generated by LLM calls in Step 2.
3.4.2 Analysis
Let us assume for notational convenience that $m_{i}=m$ for all $i\in[k]$, and $k=n/m$ is an integer.
Error metrics.
Compared to counting, there are more diverse phenomena in the sorting task in terms of error metrics. In particular, multiple possible failure modes exist in sorting with an LLM-based algorithm:
  1. 1.
    The output list might not be monotone;
  2. 2.
    The length of the output list might be larger or smaller than that of the input list;
  3. 3.
    The output list might contain numbers that do not match exactly those of the input list.
Based on these failure modes, we define the following error metrics for sorting a list, where $\bm{y}$ denotes the solution returned by the algorithm and $\bm{y}^{\star}$ denotes the ground-truth solution:
Note that the same error metrics can be similarly defined for each parallel sub-task in Step 2 of the LLM-based algorithm, and it is reasonable to expect that they become smaller as the sub-task size $m$ decreases. On the other hand, analyzing the error metrics of the overall algorithm after the final merging step can be more complicated, and might be an interesting theoretical problem on its own. As an example, focusing on the third failure mode and the $\ell_{\infty}$ error metric, we have the following guarantee, whose proof is deferred after the experiments in this subsection.
Proposition 1.
Cost metrics.
Our analysis of cost metrics for the LLM-based sorting algorithm follows Section 3.1, and is similar to that for counting. One major difference is that $\mathcal{L}_{\textsf{dec}}=O(m)$ rather than $O(1)$ for each sub-task.
@@ 3.4.2 Analysis
[…]
@@
[…]
@@
3.4.3 Experiments
We validate our analysis with numerical experiments. The input list of each problem instance is generated by randomly sampling entries of the list from the uniform distribution over the interval $[0,1]$ and then rounding each of them to two decimals.
Results with Llama-3-70B.
Our empirical results with a Llama-3-70B model are illustrated in Figure 6 and explained in the following.
@@ 3.4.3 Experiments
In Figure 6(b), we fix $n=200$ and vary the sub-task size $m$. It is confirmed that decomposing the original task into smaller parallel sub-tasks with a smaller value of $m$ implies lower error metrics achieved by the overall algorithm, while increasing certain cost metrics. Two specific observations:
@@
Figure 6: Empirical results for sorting with Llama-3-70B.
@@
Results with GPT-4-Turbo.
3.4.4 Proof of Proposition 1
@@
It is easy to check that merging multiple sorted lists is equivalent to sorting the concatenation of the lists. Therefore, we have
@@
where permute can be any permutation of $n$ elements in a list. In particular, we let permute be the permutation that sorts $[\bm{y}^{\star}_{1},\dots,\bm{y}^{\star}_{k}]$, which implies
@@
$z_{j-1}$ $-z^{\star}_{j}$ $>z_{j}$ $-z^{\star}_{j}$ $\geq$-≥ - $\epsilon,$
@@
$z_{j}$ $-z^{\star}_{j-1}$ $\geq z_{j}$ $-z^{\star}_{j}$ $\geq$-≥ - $\epsilon,$
@@
1 Input: a list $\bm{z}\in{\mathbb{R}}^{n}$ to be sorted.
@@
3.5 Example: retrieval
We study a more realistic application of LLM-based algorithms, which is answering a given question that requires retrieving some key information from a long piece of text. Our design of this task draws inspiration from the needle-in-a-haystack benchmark [Kam23] and other similar benchmarks that have been widely adopted for evaluating the long-context capability of LLMs as well as techniques of retrieval-augmented generation [WBC+15, Rou23, KBA+24, MJ23, ZCH+24].
Consider the following setting as an example. A key message (the needle) of the form “The passcode to the {targeted object, e.g. red door} is {6-digit passcode}” is randomly inserted into a piece of long text (the haystack). The algorithm is asked to answer the question “What is the passcode to the {targeted object}?”. To make the problem more challenging and fun, we let the haystack consist of alike sentences of the form “The passcode to the {colored object, e.g. red lock or green door} is {6-digit passcode}”, with colored objects different from the targeted object. This allows us to investigate both sides of retrieval capabilities of LLMs and LLM-based algorithms: retrieving the targeted information correctly, while avoiding being confused or misled by background information that might seem relevant to the question [SCM+23].
Note that while we use this concrete setting for our empirical study, the proposed algorithm and analysis in the following are actually applicable to generic settings of this retrieval task.
3.5.1 Algorithm
We consider the following LLM-based algorithm that follows the pattern of parallel decomposition:
  1. 1.
  2. 2.
    For each chunk, use one LLM call to try to answer the question based on that chunk;
  3. 3.
    Generate the final answer by majority voting.
We note a few details about this algorithm. (1) All lengths involved here are measured by the number of characters. (2) In the first step of chunking, we let each pair of adjacent chunks share an overlap that is larger than the length of the needle, to ensure that the needle will appears as a whole in at least one chunk. (3) For each LLM call in the second step, the LLM is prompted to answer “I don’t know” if it believes that there is not sufficient information in the corresponding chunk, e.g. when the chunk simply does not contain the needle. Such answers will be excluded from the final step of the algorithm. (4) In the final step of majority voting, it is possible that there are multiple (say $h$) candidate solutions with the same frequency, in which case we let the algorithm return the list of such candidates. If this list contains the ground-truth solution, we calculate the exact-match error as $1-1/h$ in our experiments.
3.5.2 Analysis
@@
Error metrics.
Let us focus on how error metrics of the LLM-based algorithm are impacted by the hyperparameter $m$. We start by identifying two failure modes of each LLM call for retrieving the targeted information from a chunk of size $m$ in Step 2 of the algorithm:
  1. 1.
    The first failure mode is that, while the needle is contained in the chunk, the LLM might mistakenly return “I don’t know” or an incorrect passcode. It is reasonable to expect that this failure mode will occur more frequently for larger values of $m$. Our early experiments with various LLMs confirmed that this failure mode starts to occur when $m$ exceeds a certain threshold specific to each LLM.
  2. 2.
    The second failure mode is that, while the needle is actually absent from the chunk, the LLM might mistakenly return a passcode that it believes is the true answer to the question. We observed empirically that this is more likely to happen when the chunk contains some objects that seem similar to the targeted object (e.g. “red lock” or “green door” vs. “red door”), and that some LLMs are more prone to this failure mode even when the value of $m$ is small, while others are less so.
Based on the above observations, we can hypothetically categorize LLMs into two types: Type-1 LLMs are only prone to the first failure mode, while Type-2 LLMs are prone to both. It turns out that analysis of error metrics for the overall LLM-based algorithm is dependent on the type of the LLM being used.
  • Analysis is simpler if a Type-1 LLM is used: a smaller value of $m$ means the first failure mode is less likely to occur in Step 2 of the algorithm, which implies higher accuracy for the final solution of the overall algorithm.
  • Analysis is more complicated if a Type-2 LLM is used, since both failure modes can possibly occur in Step 2 of the algorithm. A larger value of $m$ means the first failure mode is more likely to occur, while a smaller value of $m$ implies a larger number of chunks $k=\lceil 2n/m-1\rceil$, which can potentially increase the chance of error in the final step of majority voting, due to the frequent occurrence of the second failure mode in Step 2 of the algorithm. Consequently, the minimum error of the overall algorithm might be achieved by some intermediate value of $m$ that achieves a balance between the two failure modes. If $n$ is too large, then there might not exist a good value of $m$ that can achieve a low error, as either failure mode must occur with high probability. [^4] An informal probabilistic analysis is as follows. Given the sub-task size $m$, denote the probability of the first and second failure modes as $p_{1}(m)$ and $p_{2}(m)$ respectively. Then, the success rate of retrieval for the chunk containing the needle is $1-p_{1}(m)$, while the expected number of “false positives” from the remaining chunks is approximately $k\times p_{2}(m)\approx 2n\times p_{2}(m)/m$. One might opt for a relatively small value of $m$, which hopefully increases $1-p_{1}(m)$ and hence mitigates the first failure mode. However, even if $p_{2}(m)/m$ is very small, say $10^{-3}$, the number of false positives can still be large if the size $n$ of the original problem is large, which will cause errors in the solution returned by majority voting.
Cost metrics.
@@
Our analysis of cost metrics for this task is largely the same as that for the counting task, despite how different these two tasks might seem. Under some mild conditions explained in Section 3.1, most cost metrics of interest are monotonely decreasing in $m$, except for the end-to-end latency with parallel LLM calls, which is increasing in $m$ for parallelism degree $p=\infty$ and possibly non-monotone for finite $p$. One thing to note is that, due to the overlaps between consecutive chunks, the number of parallel sub-tasks in Step 2 of the algorithm is $k=\lceil 2n/m-1\rceil$, rather than $\lceil n/m\rceil$ as in the counting task. This implies that the value of $m$ minimizing the latency can be predicted by letting $2n/m-1\approx p$, namely $m\approx 2n/(p+1)$.
3.5.3 Experiments
@@
We validate our analysis with numerical experiments. For each task instance, the passcodes of all objects, the targeted object, the position of the haystack where the needle is inserted, etc., are all randomly chosen. The error metric of interest, namely the exact-match error, takes value 00 if the final solution of the algorithm is exactly the same as the ground-truth passcode to the targeted object, $1-1/h$ if the algorithm returns a list of $h$ candidate solutions that includes the ground-truth passcode, and $1$ otherwise.
Results with Llama-3-8B.
Figure 8 includes the results of our experiments with Llama-3-8B. [^5] After executing many LLM calls in a row during our experiments with Llama-3-8B supported by ollama [oll23], we started to observe unusually large latencies (at least two orders of magnitude larger than their normal values) for some LLM calls, even though the generated texts are normal. We believe that this is most likely due to memory-related issues caused by running ollama on a laptop with limited 16GB memory, which can be easily avoided if a laptop with more memory is used. To mitigate this issue in our experiments, we take a different approach, i.e. adding a 3-second pause between each pair of consecutive LLM calls in the LLM-based algorithm when Llama-3-8B and ollama are used. While this proves to be quite effective, anomalies might still occur after running the experiments for a long period of time, in which case we simply re-run the part of experiments containing such anomalies, or remove these data points manually before plotting if there are very few of them,
@@
Figure 8(b) includes the results for the same experiment setting, except that no needle is inserted into the haystack. One crucial observation is that the exact-match error in this case, which corresponds to the second failure mode of retrieval, remains non-zero even for very small values of $n$. This suggests that Llama-3-8B should be regarded as a Type-2 LLM prone to both failure modes.
In Figures 8(c) and 8(d), we vary the sub-task size $m$ while $n$ is fixed at $10000$ and $20000$ respectively. As was predicted by our analysis, the error of the overall algorithm is not monotone in the value of $m$, due to the presence of two failure modes. Another difference from the previous counting or sorting task is that the latency with parallelism degree $p=4$ achieves the minimum around $m=2n/(p+1)=0.4n$ rather than $m=n/p=0.25n$, which again matches our previous analysis.
Results with Llama-3-70B.
Figure 9 includes the results for the same experiments but with Llama-3-70B used within the LLM-based algorithm.
In particular, Figure 9(a) shows that Llama-3-70B achieves lower errors in the first failure mode of retrieval compared to Llama-3-8B, while Figure 9(b) suggests that Llama-3-70B is much less prone to the second failure mode and hence might be regarded as a Type-1 LLM. Consequently, in Figures 9(c) and 9(d), the exact-match error exhibits a more monotone relation with the sub-task size $m$.
Regarding the cost metrics, results with Llama-3-70B are similar to those with Llama-3-8B.
@@
Figure 8: Empirical results for retrieval with Llama-3-8B.
@@
Figure 9: Empirical results for retrieval with Llama-3-70B.
3.6 Example: retrieval-augmented generation
Our next example is a multiple-needle generalization of the previous retrieval task, which can be regarded as a synthetic and simplified version of retrieval-augmented generation (RAG) [LPP+20, GXG+24]. In particular, we consider fine-grained sentence-level retrieval by LLMs, rather than document-level or chunk-level retrieval by certain similarity measure of dense embedding vectors.

### v3

@@ A.2.2 Cost analysis
    $\mathcal{C}=\mathcal{C}(\text{sub-tasks})$ $\leq k\times\Big(\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,1)\Big)$
@@
    $\mathcal{C}=\mathcal{C}(\text{sub-tasks})$ $\leq\Big\lceil\frac{k}{p}\Big\rceil\times\Big(\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,1)\Big)$
@@ A.2.3 Experiments
Results with Llama-3-8B-Instruct.
The results are illustrated in Figure 9.
In Figure 9(a), we vary the problem size $n$ and set $m=n$ in our algorithm, in which case the algorithm becomes equivalent to a single LLM call. Unsurprising, error metrics in this counting task monotonely increase with $n$. The number of prefilling tokens increases linearly with $n$, while the number of decoding tokens is insensitive to $n$, since we prompt the LLM to output its answer directly without intermediate reasoning steps. The latency of one LLM call also increases linearly with $n$, which is likely due to the relatively small sequence lengths and the compute-bound nature of LLM inference by a Llama-3-8B-Instruct model running on a CPU.
In Figure 9(b), we fix $n=200$ and vary the hyperparameter $m$, i.e., the size of each sub-task. It is confirmed that decomposing the original task into smaller parallel sub-tasks improves the accuracy of the overall algorithm, while incuring higher cost metrics, except for the latency with infinite parallelism (which is monotonely increasing in $m$) and the latency with parallelism degree $p=4$ (which achieves the minimum at $m=n/p=50$, as was predicted by our previous analysis).
@@
Figure 10 demonstrates the empirical results for the same experiments but with GPT-4-Turbo. We observe from Figure 10(a) that the latency of one LLM call is insensitive to the input problem size $n$, which is likely because LLM inference of GPT-4-Turbo is memory-bound for the range of sequence lengths considered in our experiments. One potential implication is that the end-to-end latency with infinite parallelism $p=\infty$ might slightly increase for smaller $m$ and hence larger $k$, due to the random variation of latencies in reality; this is indeed what we observe from Figure 10(b). Other than that, the results in Figure 10 are similar to those in Figure 9.
@@
Figure 9: Empirical results for counting with Llama-3-8B-Instruct.
@@
A.3 Example: sorting
A.3.1 Algorithm details
We note two details about the sorting algorithm presented in Table 1.
  • To ensure efficiency and stability, for each LLM call, we prompt the LLM to generate the sorted list directly, without intermediate reasoning steps. It has been verified empirically that the LLMs considered in our experiments can follow such instructions, and generate text of length $\mathcal{L}_{\textsf{dec}}(m)=O(m)$ that can be easily parsed into a list.
  • Step 3 of the algorithm relies on a classical symbolic algorithm for merging two sorted lists, which maintains two moving pointers, one for each list, and chooses each entry of the merged list by comparing the values corresponding to the two pointers. Merging multiple sorted lists can be done by merging one pair of lists at a time, in an incremental or hierarchical manner. Python code for these procedures can be found in Listing 1 at the end of this subsection. Although they are designed under the assumption that the input lists are sorted, they can also be applied to input lists that are not fully sorted, which can possibly happen within the LLM-based algorithm since the input lists in Step 3 are generated by LLM calls in Step 2.
A.3.2 Cost analysis
Cost analysis of for the LLM-based sorting algorithm follows Section 3.1, and is similar to that for counting. One major difference is that $\mathcal{L}_{\textsf{dec}}=O(m)$ rather than $O(1)$ for each sub-task.
@@ A.3.2 Cost analysis
    $\mathcal{C}=\mathcal{C}(\text{sub-tasks})$ $\leq k\times\Big(\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,m)\Big)$
@@
    $\mathcal{C}=\mathcal{C}(\text{sub-tasks})$ $\leq\Big\lceil\frac{k}{p}\Big\rceil\times\Big(\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,m)\Big)$
@@
A.3.3 Experiments
In our experiments, the input list of each problem instance is generated by randomly sampling entries of the list from the uniform distribution over the interval $[0,1]$ and rounding each of them to two decimals.
Results with Llama-3-70B-Instruct.
The results are illustrated in Figure 11.
@@ A.3.3 Experiments
In Figure 11(b), we fix $n=200$ and vary the sub-task size $m$. It is confirmed that decomposing the original task into smaller parallel sub-tasks implies lower error metrics achieved by the overall algorithm, while increasing certain cost metrics. Two specific observations:
@@
Figure 11: Empirical results for sorting with Llama-3-70B-Instruct.
@@
A.3.4 Proof of Proposition 1
@@
To start with, it is easy to check that merging multiple sorted lists is equivalent to sorting the concatenation of the lists. Therefore, we have
@@
where permute can be any permutation of $n$ elements in a list. In particular, we let permute be the permutation that sorts $\bm{y}^{\star}_{1},\dots,\bm{y}^{\star}_{k}]$, which implies
@@
$z_{j-1}$ $-z^{\star}_{j}$ $>z_{j}$ $-z^{\star}_{j}$ $\geq$- $\epsilon,$
@@
$z_{j}$ $-z^{\star}_{j-1}$ $\geq z_{j}$ $-z^{\star}_{j}$ $\geq$- $\epsilon,$
@@
1Input: a list $\bm{z}\in{\mathbb{R}}^{n}$ to be sorted.
@@
A.4 Example: retrieval
Our investigation of the retrieval task draws inspiration from the needle-in-a-haystack benchmark (Kamradt, 2023) and other similar benchmarks that have been widely adopted for evaluating the long-context capability of LLMs as well as techniques of retrieval-augmented generation (Weston et al., 2015; Roucher, 2023; Kuratov et al., 2024; Mohtashami & Jaggi, 2023; Zhang et al., 2024a). For concreteness, consider the following setting. A key message (the needle) of the form “The passcode to the {targeted object, e.g., red door} is {6-digit passcode}” is randomly inserted into a piece of long text (the haystack). The algorithm is asked to answer the question “What is the passcode to the {targeted object}?”. To make the problem more challenging and fun, we let the haystack consist of alike sentences of the form “The passcode to the {colored object, e.g., red lock or green door} is {6-digit passcode}”, with colored objects different from the targeted object. This allows us to investigate both sides of retrieval capabilities of LLMs and LLM-based algorithms: retrieving the targeted information correctly, while avoiding being confused or misled by background information that might seem relevant to the question (Shi et al., 2023). Note that while we use this concrete setting for our empirical study, the proposed algorithm and analysis in the following are actually applicable to generic cases of this retrieval task.
A.4.1 Algorithm details
We note a few details about the LLM-based retrieval algorithm presented in Table 1.
  • All lengths involved in the algorithm are measured by the number of characters.
  • In the first step of chunking, we let each pair of adjacent chunks share an overlap that is larger than the length of the needle, to ensure that the needle will appears as a whole in at least one chunk.
  • For each LLM call in the second step, the LLM is prompted to answer “I don’t know” if it believes that there is not sufficient information in the corresponding chunk, e.g., when the chunk simply does not contain the needle. Such answers will be excluded from the final step of the algorithm.
  • In the final step, i.e., majority voting, it is possible that there exist multiple (say $h$) candidate solutions with the same frequency, in which case we let the algorithm return the list of such candidates. If this list contains the ground-truth solution, we calculate the exact-match error as 11/h1-1/h in our experiments.
A.4.2 Cost analysis
@@
Our cost analysis for this task is largely the same as that for the counting task, despite how different these two tasks might seem. Under some mild conditions explained in Section 3.1, most cost metrics of interest are monotonely decreasing in $m$, except for the end-to-end latency with parallel LLM calls, which is increasing in $m$ for parallelism degree $p=\infty$ and possibly non-monotone for finite $p$. One thing to note is that, due to the overlaps between consecutive chunks, the number of parallel sub-tasks in Step 2 of the algorithm is $k=\lceil 2n/m-1\rceil$, rather than $\lceil n/m\rceil$ as in the counting task. This implies that the value of $m$ minimizing the latency can be predicted by letting $2n/m-1\approx p$, namely $m\approx 2n/(p+1)$.
A.4.3 Experiments
@@
We validate our analysis with numerical experiments. For each task instance, the passcodes of all objects, the targeted object, the position of the haystack where the needle is inserted, etc., are all randomly chosen. The error metric of interest, namely the exact-match error, takes value 0 if the final solution of the algorithm is exactly the same as the ground-truth passcode to the targeted object, 11/h1-1/h if the algorithm returns a list of $h$ candidate solutions that includes the ground-truth passcode, and $1$ otherwise.
Results with Llama-3-8B-Instruct.
The results are illustrated in Figure 13.
@@
Figure 13(b) includes the results for the same experiment setting, except that no needle is inserted into the haystack. One crucial observation is that the exact-match error in this case, which corresponds to the second failure mode of retrieval, remains non-zero even for very small values of $n$. This suggests that Llama-3-8B-Instruct should be regarded as a Type-B LLM that is prone to both failure modes explained in Section 3.2.3.
In Figures 13(c) and 13(d), we vary the sub-task size $m$ while $n$ is fixed at $10000$ and $20000$ respectively. As was predicted by our analysis, the error of the overall algorithm is not monotone in the value of $m$, due to the presence of two failure modes. Another difference from the previous counting or sorting task is that the latency with parallelism degree $p=4$ achieves the minimum around $m=2n/(p+1)=0.4n$ rather than $m=n/p=0.25n$, which again matches our previous analysis.
Results with Llama-3-70B-Instruct.
Figure 14 includes the results for the same experiments but with Llama-3-70B-Instruct used within the LLM-based algorithm. In particular, Figure 14(a) shows that this LLM achieves lower errors in the first failure mode of retrieval compared to Llama-3-8B-Instruct, while Figure 14(b) suggests that it is much less prone to the second failure mode and hence might be regarded as a Type-A LLM. Consequently, in Figures 14(c) and 14(d), the exact-match error exhibits a more monotone relation with the sub-task size $m$. Regarding the cost metrics, results with Llama-3-70B-Instruct are similar to those with Llama-3-8B-Instruct.
@@
Figure 13: Empirical results for retrieval with Llama-3-8B-Instruct.
@@
Figure 14: Empirical results for retrieval with Llama-3-70B-Instruct.
A.5 Example: retrieval-augmented generation
Our next example is a multiple-needle generalization of the previous retrieval task, which can be regarded as a synthetic and simplified version of retrieval-augmented generation (RAG) (Lewis et al., 2020; Gao et al., 2024b). In particular, we consider fine-grained sentence-level retrieval by LLMs, rather than document-level or chunk-level retrieval by certain similarity measure of dense embedding vectors.